OpenStudy (anonymous):

(-2a^2)^-2 ____________ 4a^2

OpenStudy (anonymous):

\[\frac{(-2a^2)^{-2}}{4a^2}=\frac{\frac{1}{4a^4}}{4a^2}=\frac{1}{16a^6}\]

OpenStudy (anonymous):

Did you add a^4 and a^2 to get a^6 at the bottom?

OpenStudy (anonymous):

Multiplied, which causes the exponents to be added.